OpenStudy (anonymous):

@jim_thompson5910 Math is a pain..

OpenStudy (anonymous):

It's confusing me..

jimthompson5910 (jim_thompson5910):

( h o h)(10) is the same as h( h(10) )

jimthompson5910 (jim_thompson5910):

h(10) = ??

OpenStudy (anonymous):

so I would do 6-x (6-x (10) ) h= 6-x so h(10) means 6-x(10) ? If so, that means 6-x (10)= 6-10x

jimthompson5910 (jim_thompson5910):

h(x) = 6 - x h(10) = 6 - 10 h(10) = ???

OpenStudy (anonymous):

= -4 ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

h(10) = -4

jimthompson5910 (jim_thompson5910):

h( h(10) ) = h( -4 ) = ???

OpenStudy (anonymous):

so confused @jim_thompson5910 ...

jimthompson5910 (jim_thompson5910):

you saw how I got h(10) right?

OpenStudy (anonymous):

@jim_thompson5910 It's a function of a function, a composite function, the right answer is 10

OpenStudy (anonymous):

\[(h\cdot h)=h[h(x)]\]if\[h(x)=6-x\]then\[h[h(x)]=6-h(x)\]replacing x by 10\[h(x)=6-10=-4\]replacing this new value at the composite function\[h[h(10)]=6-h(10)=6-(-4)=\boxed{10}\]therefore\[\boxed{\boxed{(h\cdot h)=10}}\]

OpenStudy (anonymous):

Thank you! ^

OpenStudy (anonymous):

really Great Explanation ! :DD

OpenStudy (anonymous):

You're welcome ;)